Given 2A (g) + B (g) 3C (g) + D (g)

Initially 6 moles of A and 10 moles of B enclosed in 2 litre container at equilibrium 30% of A has reacted. Find Kc.

Grade:11

1 Answers

Shivank Kapila
6 years ago
See we have the reaction 2A + B -------> 3C + DAt t=0. 6. 10. - -At equil. 4.2. 9.1. 5.4. 1.8 becauseIt is written that 30% A has reactedNow, Kc = [C]^3 [D]/[A]^2[B]Dividing the moles written above by 2 because we need conc. Conc. Of A = 2.1, B= 9.1/2, C= 5.4/2, D= 0.9Put values in Kc and you get the answerHope it helps......